The table below shows the number of state-registered automatic weapons and the murder rate for several Northwestern states.

xx 11.6 8.6 7.1 3.3 2.5 2.7 2.5 0.9
yy 13.8 11.3 10.2 6.6 5.9 6.5 6.5 4.9

xx = thousands of automatic weapons
yy = murders per 100,000 residents

This data can be modeled by the equation y=0.84x+4.1.y=0.84x+4.1. Use this equation to answer the following;
Special Note: I suggest you verify this equation by performing linear regression on your calculator.

A) How many murders per 100,000 residents can be expected in a state with 6.1 thousand automatic weapons?

Answer =  Round to 3 decimal places.

B) How many murders per 100,000 residents can be expected in a state with 4 thousand automatic weapons?

Answer =  Round to 3 decimal places.
Answer #1

Answer:

Given that

X :  thousands of automatic weapons
Y = murders per 100,000 residents

modelled equation:

y=0.84x+4.1

A)

murderedr per 100,000 residents with 6.1 thousand automatic weapons

x=6.1

y=0.84x+4.1
=0.84(6.1)+4.1

y=9.224

9.224 murders per 100,000 residents can be expected in a state with 6.1 thosands automatic weapons.

B)

murders per 100,000 residents with 4 thousand automatic weapons

x=4

y=0.84(4)+4.1

=3.36+4.1
y=7.46.